# SimuRA: Towards General Goal-Oriented Agent via Simulative Reasoning Architecture with LLM-Based World Model

### v1

@@ Abstract
AI agents built on large language models (LLMs) hold enormous promise, but current practice focuses on a one-task-one-agent approach, which not only falls short of scalability and generality, but also suffers from the fundamental limitations of autoregressive LLMs. On the other hand, humans are general agents who reason by mentally simulating the outcomes of their actions and plans. Moving towards a more general and powerful AI agent, we introduce SimuRA, a goal-oriented architecture for generalized agentic reasoning. Based on a principled formulation of optimal agent in any environment, SimuRA overcomes the limitations of autoregressive reasoning by introducing a world model for planning via simulation. The generalized world model is implemented using LLM, which can flexibly plan in a wide range of environments using the concept-rich latent space of natural language. Experiments on difficult web browsing tasks show that SimuRA improves the success of flight search from 0% to 32.2%. World-model-based planning, in particular, shows consistent advantage of up to 124% over autoregressive planning, demonstrating the advantage of world model simulation as a reasoning paradigm. We are excited about the possibility for training a single, general agent model based on LLMs that can act superintelligently in all environments. To start, we make ReasonerAgent-Web, a web-browsing agent built on SimuRA with pretrained LLMs, available as a research demo for public testing.
@@ 1 Introduction
AI agents powered by large language models (LLMs) hold tremendous potential for handling tasks that require flexible decision making. Recently, there have been great advancements in agents specialized in web and computer automation openai_cua_2025 ; deepmind2024mariner ; zhou2023webarena ; xie2024osworldbenchmarkingmultimodalagents , internet research openai_deepresearch_2025 ; anthropic2024claude35sonnet ; google_gemini_deep_research_2024 , social simulation park2023generativeagentsinteractivesimulacra , software development anysphere2025cursor ; wang2025openhandsopenplatformai , scientific research gottweis2025towards ; yamada2025aiscientistv2workshoplevelautomated , and so on. Despite the promise, current LLMs often prove insufficient for solving complex agentic tasks, suffering from issues such as hallucination, repetitions, or failure at complex planning liu2025advanceschallengesfoundationagents ; chakraborty2025healempiricalstudyhallucinations . To address these issues, many approaches focus on creating agents tailored to specific tasks like the above examples. However, this strategy have some inherent drawbacks. Economically, redesigning custom agents for every task is costly and not scalable from a business standpoint. Intellectually, narrowly focused solutions offer no clear path towards general and transferrable intelligence. LU2020221 Technically, autoregressive LLMs rely on linear, step-by-step reasoning that often leads to errors that propagate through their thought trajectories andreas2022languagemodelsagentmodels ; yao2023react . Humans, in contrast, are generalist problem-solvers that can reason and plan to achieve goals in diverse environments. Using a single cognitive system, we adapt to different tasks not only by linear reasoning, but also by imagining potential outcomes, simulating possibilities using a mental world model, and planning accordingly. lecun2022pathtowards
@@
Figure 1: Demo of tasks performed using a web-browsing agent built on SimuRA with simulative planning using a LLM-based world model.
@@
Moving towards a more general and powerful AI agent, we introduce SimuRA (Simulative Reasoning Architecture), a goal-oriented architecture for generalized agentic reasoning. SimuRA mitigates the limitations of LLM autoregressive reasoning by introducing world model as the engine for planning via simulation. Specifically, a policy module first proposes a few potential actions, aimed at achieving specific goals based on agent identity and environment. Then, the world model simulates the outcomes of those proposed actions. Finally, a critic module evaluates these outcomes against the initial goals in order to select the best action from the candidates. Because simulating the full details of the world is infeasible and unnecessary for planning, we extract only the relevant information using natural language as a compact but complete representation, and simulate the next world in this latent space. To ensure robustness from observation noise and distracting execution details, we further propose a hierarchical architecture that isolates perception, simulative planning, and action selection which ensures adaptability and consistency across diverse tasks. Experiments on a range of web browsing tasks show SimuRA improving substantially compared to baselines, increasing the success rate of flight search from 0% to 32.2%, with reasoning by WM simulation outperforming LLM autoregressive reasoning by up to 124%. Figure 1 shows examples of the agent performing multi-website, long-range task such as flight searching, online shopping, and news research.
@@
For evaluation and demonstration purposes, we implemented SimuRA as an open-source library available via LLM Reasoners hao2024llm . [^1] https://github.com/maitrix-org/llm-reasoners/tree/main/examples/ReasonerAgent-Web The resulting web agent, ReasonerAgent-Web reasoneragent2025 , is available as a research preview. [^2] https://easyweb.maitrix.org/ We are actively expanding the system to address broader challenges and to further demonstrate its generality across a wider range of task domains.
@@ 2 Related Work
Model-based planning for agents have long been frequently discussed and studied. Early work demonstrated the success of this approach by testing in classic games like go, chess, shogi and Atari. oh2015actionconditionalvideopredictionusing ; Schrittwieser_2020 . Later on, world model was used for policy optimization and experimented on control tasks. janner2021trustmodelmodelbasedpolicy ; hansen2022temporaldifferencelearningmodel In recent years, with the boost in foundation model’s capabilities, world-model-based planning was applied to more complex problems like math reasoning hao2023reasoninglanguagemodelplanning , playing Minecraft hafner2024masteringdiversedomainsworld , and web browsing gu2025llmsecretlyworldmodel . However, these world models typically represent and predict the world states using holistic continuous embeddings, which suffer from noise and high variability which detracts from robust and stable decision-making barrett2017emotions . SimuRA instead adopts natural language as a discrete, concept-based latent space for consistent representation and prediction, which shows more general applicability across tasks in practice.
@@
Web browsing and navigation were chosen to evaluate SimuRA due to their realism and the complex decision-making they demand across diverse, dynamic interfaces. Recent years have seen the emergence of several prominent web-browsing agents, from proprietary ones such as OpenAI’s Operator openai_cua_2025 , Anthropic’s Computer Use anthropic2024claude35sonnet , and Google-DeepMind’s Project Mariner deepmind2024mariner , and open-source ones including OpenHand’s BrowsingAgent openhands , WebVoyager he2024webvoyagerbuildingendtoendweb , CogAgent hong2024cogagentvisuallanguagemodel and WebAgent gur2024realworldwebagentplanninglong . These agents are typically built on simple ReAct-based yao2023react autoregressive reasoning which have difficulty recovering from previous mistakes; their often specialized design also preclude these approaches from generalizing to other task domains like social interactions and the physical world. Numerous benchmarks have been introduced to evaluate these web agents, including WebArena zhou2023webarena , WebVoyager he2024webvoyagerbuildingendtoendweb MiniWoB++ liu2018reinforcement , Mind2Web deng2023mind2webgeneralistagentweb , and WebShop yao2023webshopscalablerealworldweb . Despite wide adoption, these benchmarks are usually either built in simulated and simplified environments, based on outdated questions, or lacks convincing method of measuring task completion, which detract from the goal of evaluating practically useful web agents. To address these challenges, we build FlightQA, an new dataset for evaluating agent ability in real-time complex website navigation. More details are included in Section 4.1.
@@
There have been various attempts of building generalist agents recently. One major approach focuses on creating a multi-agent system that consists of a unified interface on top of a few specialist agents that collaborates to decompose and complete complex tasks. hu2025owloptimizedworkforcelearning ; jiang2025longtermmemoryfoundation ; zhang2025agentorchestrahierarchicalmultiagentframework ; fourney2024magenticonegeneralistmultiagentsolving Although this approach could lead to impressive performance on benchmarks, it has a few inherent limitations. First of all, tasks in reality could be versatile and may constantly require new specialist agents to be added to the system to achieve optimal performance, which is not efficient. Moreover, independently trained specialist agents for different domains are unable to leverage shared experience in the way that world model training enables. Finally, error propagation along the interaction trajectory remains an open challenge and is further complicated by the presence of multiple agents. Another popular approach utilizes framework similar to the CodeActAgent wang2024executablecodeactionselicit . These agents openhands ; smolagents ; qiu2025alitageneralistagentenabling suffer from inaccurate code plans and have limited ability to revise or correct prior errors as well. SimuRA, on the other hand, is able to avoid these limitations by working as a monolithic architecture in which world model act as a central planning component.
@@ 3.1 Formulation of Agent-Environment Model
We first present our formulation of an optimal goal-oriented agent following the agent-environment model presented in xing2025critiques : We consider an agent $\pi$ with identity $i$ (e.g., name, description) and goal $g$ acting in environment $\mu$ (e.g., web browser, physical world, the entire universe) with action space $\mathcal{A}$ and state space $\mathcal{S}$. Formally, at each time step $t$, the agent $\pi$ takes the current state $s_{t}\in\mathcal{S}$ and outputs the next action $a_{t}\in\mathcal{A}$ following a policy distribution $p_{\pi}(a_{t}\mid s_{t})$, while the environment $\mu$ takes the current state $s_{t}$ and action $a_{t}$, and outputs the next state $s_{t+1}\in\mathcal{S}$ based on the distribution $p_{\mu}(s_{t+1}|s_{t},a_{t})$. We can thus denote the distribution of the interaction trajectory up to timestep $T$, or $(a_{t},s_{t+1},\dots,a_{T-1},s_{T})$ given the current state $s_{t}$ as below:
@@
$V_{\pi,\mu}^{g}(s_{t})$ :$\vcentcolon=\mathbb{E}_{\pi,\mu}\left[\sum_{k=t}^{\infty}\gamma_{k}r(g,s_{k})\mathrel{\bigg{|}}s_{t}\right]$
@@ 3.2 Definition of Optimal Agent
$\pi^{*}_{\mu}(s_{t})={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}}}_{\text{possible actions}}}\ \sum_{s_{t+1:T}}\Bigg{(}{\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,s_{k})+\gamma_{T}V_{\pi,\mu}^{g}(s_{T})}_{\text{goal progress}}}\Bigg{)}\prod_{i=t}^{T-1}\ {\underbrace{p_{\mu}(s_{i+1}\mid s_{i},a_{i})}_{\text{universe response}}}$ (4)
In practice, agents often samples promising action candidates using a policy function $\tilde{\pi}$ through the distribution $p_{\tilde{\pi}}(a_{t}\mid s_{t})$. Building the optimal agent thus requires capabilities for proposing possible actions ($\tilde{\pi}$), predicting their outcomes ($\mu$), and evaluating goal progress ($r,V$), respectively. Note that typical reactive agents that output the next action directly can be seen as taking the first sample from $\tilde{\pi}$ (similar to “System 1” in humans which makes fast, instinctive reactions kahneman2011thinking ), without simulating and evaluating the outcomes using $\mu$ and $V$ (similar to “System 2” responsible for deliberate decision-making). In terms of LLM-based agents, this can also be seen as the agent generating a plan using autoregressive LLMs, which has no way of correcting errors during the sampling process.
@@ 3.3 World Model for Generalized Simulative Reasoning
Note that the optimal decision-making process defined in Equation 4 requires the agent to have access to the ground-truth world state $s$ and the environment $\mu$ to experience and optimize over. However, these are often not available aside from simple scenarios like Go and Chess games (silver2016mastering, ; silver2017mastering, ) – imagine building an spacecraft to land on Mars, or simply a humanoid robot relying on noisy sensors in daily environments. World Model (WM) thus arises as a crucial component for predicting any environment’s response to a general agent. Specifically, a WM $f$ operates on an internal representation of the world state, denoted as a belief state $\hat{s}_{t}$, which is derived from sensory inputs $o_{t}$ via an Encoder $h$ (unlike the optimal agent described in §3.2 which has direct access to the true world state $s_{t}$). Given proposed action $a_{t}$, the WM predicts the next belief state $\hat{s}_{t+1}$ according to the distribution $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a_{t})$. The predicted belief state then allows the agent to propose the next action, continuing the cycle of prediction and action up to the desired time horizon $T$. Thus, a WM here essentially functions as a generative model of possible future world states, which enables simulative reasoning, or “thought experiments". Formally, for the optimal agent $\pi^{*}_{f}$ equipped with WM $f$ in belief state $\hat{s}_{t}$, we define the simulation-based decision rule in Equation 6 as follows:
[…]
@@
3.4 Design of Simulative Reasoning Agent Using LLM-Based World Model
In this subsection, we present our design of a generalizable simulative reasoning agent using large language models (LLMs) as building blocks due to the latter’s strength in a wide range of capabilities such as summarization, commonsense knowledge, reflection, and tool use, which are gained from large-scale pretraining and instruction tuning. In particular, we provide detailed discussion on design decisions that enable robust and general applicability across environments and tasks.

### v2

# SimuRA: A World-Model-Driven Simulative Reasoning Architecture for General Goal-Oriented Agents

@@ Abstract
AI agents built on foundation models hold enormous promise. Current practice, however, focuses on a one-task-one-agent approach, which not only falls short of scalability and generality, but also faces practical limitations from black-box autoregressive reasoning, where decisions unfold token by token without explicit simulation or counterfactual evaluation of outcomes. Humans, on the other hand, reason and plan by mentally simulating the consequences of actions within an internal model of the world – a capability that supports flexible, goal-directed behavior across diverse contexts. Moving towards a more general and powerful AI agent, we introduce SimuRA, a goal-oriented architecture for generalized agentic reasoning. Based on a principled formulation of an optimal agent in any general environment, SimuRA addresses the limitations of black-box autoregressive reasoning by incorporating the world model for planning via simulation. Our prototype world model is implemented using LLMs as a substrate, leveraging the natural language as a discrete, hierarchical representation grounded in concepts for planning, while remaining model-agnostic. On complex web-browsing tasks such as flight search, SimuRA improves the success rate from 0% to 32.2% compared to a representative open-web agent baseline. Across tasks, world-model-based planning achieves up to 124% higher task completion rates than a matched black-box autoregressive baseline, demonstrating the advantages of simulative reasoning. [^1] All experiments were conducted in 2024 with then-available models and tooling. We release ReasonerAgent-Web, a web-browsing agent built on SimuRA, as an open-source research demo.
@@ 1 Introduction
AI agents powered by foundation models such as Large Language Models (LLM) or Vision Language Models (VLM) have demonstrated tremendous potential for handling tasks that require flexible decision making. For instance, there have been releases of agents specialized in web and computer automation openai_cua_2025 ; deepmind2024mariner ; zhou2023webarena ; xie2024osworldbenchmarkingmultimodalagents , internet research openai_deepresearch_2025 ; anthropic2024claude35sonnet ; google_gemini_deep_research_2024 , social simulation park2023generativeagentsinteractivesimulacra , software development anysphere2025cursor ; wang2025openhandsopenplatformai , scientific research gottweis2025towards ; yamada2025aiscientistv2workshoplevelautomated , and so on. Despite the promise, these agent architectures remain tailored to specific tasks, limiting their scalability and transferability. Recent reasoning LLMs optimized via end-to-end reinforcement learning (RL) openai_learning_to_reason ; guo2025deepseek have begun to show traces of emergent planning and improved long-horizon control, suggesting that black-box autoregressive reasoning can partially internalize aspects of deliberative reasoning.  [^2] In this work, we use “autoregressive reasoning” to denote the procedural mode of decision-making that iteratively predicts the next abstract internal state $z_{t}$ (e.g., text or latent tokens) based on distribution $p(z_{t}\mid z_{<t},x)$ given input $x$, without explicit modeling or simulation of future outcomes – distinct from autoregressive modeling as a statistical or architectural property. However, their reasoning still proceeds primarily through token-by-token generation, lacking an explicit mechanism for simulating and evaluating counterfactual futures. As a result, even strong models can remain vulnerable to locally myopic decisions or error accumulation over extended trajectories andreas2022languagemodelsagentmodels ; yao2023react . Humans, in contrast, are capable of reasoning and planning to achieve goals in diverse environments. Using a single cognitive system, human beings adapt to different task and environments not only by linear reasoning, but also by imagining potential outcomes, simulating possibilities using a mental world model, and planning accordingly Ball_2020 ; lecun_path_2022 .
@@
Figure 1: Demo of tasks performed using a web-browsing agent built on SimuRA with simulative planning using a world model.
@@
Moving towards more powerful and generally applicable AI agents, we introduce SimuRA (Simulative Reasoning Architecture), a goal-oriented architecture for generalized agentic reasoning. SimuRA mitigates the limitations of linear autoregressive reasoning by introducing world model as the engine for planning via simulation. Specifically, a policy module first proposes a few potential actions, aimed at achieving specific goals based on agent identity and environment. The world model then simulates the outcomes of those proposed actions. Finally, a critic module evaluates these outcomes against the initial goals in order to select the best action from the candidates. Because simulating the full details of the world is infeasible and unnecessary for planning, we extract only the relevant information using natural language as a compact, semantically structured representation space for reasoning. To ensure robustness to observation noise and distracting execution details, we further propose a hierarchical architecture that isolates perception, simulative planning, and action selection, enhancing adaptability and consistency across diverse tasks.
We conducted experiments by implementing these modular components within a LLM-based pipeline, applied to a range of web browsing tasks such as complex website navigation, multi-hop, multi-website QA, and general web automation. In particular, to evaluate the ability of agents in robustly interacting with complex websites, we develop FlightQA (Section 4.1), a dataset of flight search questions with varying number of constraints and types of constraints (e.g., dates, transfer, pricing) that are time-insensitive and automatically verifiable. Results show that SimuRA performs substantially better compared to baselines, increasing the success rate on FlightQA from 0% to 32.2% compared to a popular open-web agent. In particular, world-model-based planning outperforms a matched autoregressive reasoning baseline by up to 124% in task-completion rate, showing the advantage of simulative reasoning. All experiments were conducted in 2024 with the then-available models and browser tooling. Figure 1 shows examples of the agent performing multi-website, long-range task such as flight searching, online shopping, and news research.
@@
For evaluation and demonstration purposes, we implemented SimuRA as an open-source library available via LLM-Reasoners hao2024llm . [^3] https://github.com/maitrix-org/llm-reasoners/tree/main/examples/ReasonerAgent-Web The resulting web agent, ReasonerAgent-Web reasoneragent2025 , is available as an open-source research demo. [^4] https://easyweb.maitrix.org/ We are actively developing next versions of the system to address broader challenges and to further demonstrate its generality across a wider range of tasks and environments.
@@ 2 Related Work
Model-based planning for agents have long been frequently discussed and studied. Early work demonstrated the success of this approach by testing in classic games like go, chess, shogi and Atari. oh2015actionconditionalvideopredictionusing ; Schrittwieser_2020 . Later on, world model was used for policy optimization and experimented on control tasks. janner2021trustmodelmodelbasedpolicy ; hansen2022temporaldifferencelearningmodel In recent years, with the boost in foundation model’s capabilities, world-model-based planning was applied to more complex problems like math reasoning hao2023reasoninglanguagemodelplanning , playing Minecraft hafner2024masteringdiversedomainsworld , and web browsing gu2025llmsecretlyworldmodel . However, these world models typically represent and/or predict the world states using holistic continuous embeddings, which suffer from noise and high variability which detracts from robust and stable decision-making barrett2017emotions . SimuRA instead adopts natural language as a discrete, concept-based latent space for consistent representation and prediction, which shows more general applicability across tasks in practice.
@@
Web browsing and navigation were chosen to evaluate SimuRA due to their realism and the complex decision-making they demand across diverse, dynamic interfaces. Recent years have seen the emergence of several prominent web-browsing agents, from proprietary ones such as OpenAI’s Operator openai_cua_2025 , Anthropic’s Computer Use anthropic2024claude35sonnet , and Google-DeepMind’s Project Mariner deepmind2024mariner , to open-source ones including OpenHand’s BrowsingAgent openhands , WebVoyager he2024webvoyagerbuildingendtoendweb , CogAgent hong2024cogagentvisuallanguagemodel and WebAgent gur2024realworldwebagentplanninglong . These agents are typically built on simple ReAct-based yao2023react autoregressive reasoning which can have difficulty recovering from previous mistakes; their often specialized design also preclude these approaches from generalizing to other task domains like social interactions and the physical world. Numerous benchmarks have been introduced to evaluate these web agents, including WebArena zhou2023webarena , WebVoyager he2024webvoyagerbuildingendtoendweb MiniWoB++ liu2018reinforcement , Mind2Web deng2023mind2webgeneralistagentweb , and WebShop yao2023webshopscalablerealworldweb . Despite wide adoption, these benchmarks are usually either built in simulated and simplified environments, based on outdated questions, or lacks convincing method of measuring task completion, which detract from the goal of evaluating practically useful web agents. To address these challenges, we build FlightQA, an new dataset for evaluating agent ability in real-time complex website navigation. More details are included in Section 4.1.
@@
There have been various attempts of building generalist agents recently. One class of methods focuses on building multi-agent workflows where different tasks or subtasks are delegated to different specialist agents hu2025owloptimizedworkforcelearning ; jiang2025longtermmemoryfoundation ; zhang2025agentorchestrahierarchicalmultiagentframework ; fourney2024magenticonegeneralistmultiagentsolving . While versatile and extensible, improvements of the system often requires implementation of new specialist agents. We aim to build a holistic architecture capable of leveraging shared experience across different domains and modalities, a core benefit enabled by world model learning. Alternatively, there is another major line of work focus on composing executable scripts to flexibly control predefined tools, thus enabling generalist behavior wang2024executablecodeactionselicit ; openhands ; smolagents ; qiu2025alitageneralistagentenabling . Our goal is to further enhance the flexibility by allowing the agent to make high-level plans with natural language, think ahead, update its plans, and correct potential mistakes during execution.
@@ 3.1 Formulation of Agent-Environment Model
We first present our formulation of an optimal goal-oriented agent following the agent-environment model presented in xing2025critiques : We consider an agent $\pi$ with identity $i$ (e.g., name, description) and goal $g$ acting in environment $\mu$ (e.g., web browser, physical world, the entire universe) with action space $\mathcal{A}$ and state space $\mathcal{S}$ (Figure 2). Formally, at each time step $t$, the agent $\pi$ takes the current state $s_{t}\in\mathcal{S}$ and outputs the next action $a_{t}\in\mathcal{A}$ following a policy distribution $p_{\pi}(a_{t}\mid s_{t})$, while the environment $\mu$ takes the current state $s_{t}$ and action $a_{t}$, and outputs the next state $s_{t+1}\in\mathcal{S}$ based on the distribution $p_{\mu}(s_{t+1}|s_{t},a_{t})$. We can thus denote the distribution of the interaction trajectory up to timestep $T$, or $(a_{t},s_{t+1},\dots,a_{T-1},s_{T})$ given the current state $s_{t}$, as below:
@@
$V_{\pi,\mu}^{g}(s_{t})$ $\vcentcolon=\mathbb{E}_{\pi,\mu}\left[\sum_{k=t}^{\infty}\gamma_{k}r(g,s_{k})\mathrel{\bigg|}s_{t}\right]$
@@ 3.2 Definition of Optimal Agent
$\pi^{*}_{\mu}(s_{t})={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}}}_{\text{possible actions}}}\ \sum_{s_{t+1:T}}\Bigg({\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,s_{k})+\gamma_{T}V_{\pi,\mu}^{g}(s_{T})}_{\text{goal progress}}}\Bigg)\prod_{i=t}^{T-1}\ {\underbrace{p_{\mu}(s_{i+1}\mid s_{i},a_{i})}_{\text{universe response}}}\$(4)
In practice, agents often samples promising action candidates using a policy function $\tilde{\pi}$ through the distribution $p_{\tilde{\pi}}(a_{t}\mid s_{t})$. Building the optimal agent thus requires capabilities for proposing possible actions ($\tilde{\pi}$), predicting their outcomes ($\mu$), and evaluating goal progress ($r,V$), respectively. Note that typical reactive agents that output the next action directly can be seen as taking the first sample from $\tilde{\pi}$ (similar to “System 1” in humans which makes fast, instinctive reactions kahneman2011thinking ), without simulating and evaluating the outcomes using $\mu$ and $V$ (similar to “System 2” responsible for deliberate decision-making), which provide opportunities for spotting and correcting errors from the deliberation process.
@@ 3.3 World Model for Generalized Simulative Reasoning
Note that the optimal decision-making process defined in Equation 4 requires the agent to have access to the ground-truth world state $s$ and the environment $\mu$ to experience and optimize over. However, these are often not available aside from simple scenarios like Go and Chess games (silver2016mastering, ; silver2017mastering, ) – imagine building a spacecraft to land on Mars, or simply a humanoid robot relying on noisy sensors in daily environments. World Model (WM) thus arises as a crucial component for predicting any environment’s response to a general agent. Specifically, a WM $f$ operates on an internal representation of the world state, denoted as a belief state $\hat{s}_{t}$, which is derived from sensory inputs $o_{t}$ via an Encoder $h$ (unlike the optimal agent described in §3.2 which has direct access to the true world state $s_{t}$). Given proposed action $a_{t}$, the WM predicts the next belief state $\hat{s}_{t+1}$ according to the distribution $p_{f}(\hat{s}_{t+1}\mid\hat{s}_{t},a_{t})$. The predicted belief state then allows the agent to propose the next action, continuing the cycle of prediction and action up to the desired time horizon $T$. Thus, a WM here essentially functions as a generative model of possible future world states, which enables simulative reasoning, or “thought experiments". Formally, for the optimal agent $\pi^{*}_{f}$ equipped with WM $f$ in belief state $\hat{s}_{t}$, we define the simulation-based decision rule in Equation 6 as follows:
$\pi^{*}_{f}(\hat{s}_{t})={\underbrace{\operatorname*{arg\,max}_{a_{t:T-1}}}_{\text{possible actions}}}\ \sum_{\hat{s}_{t+1:T}}\Bigg({\underbrace{\sum_{k=t}^{T-1}\gamma_{k}r(g,\hat{s}_{k})+\gamma_{T}V_{\pi,f}^{g}(\hat{s}_{T})}_{\text{goal progress}}}\Bigg)\prod_{i=t}^{T-1}\ {\underbrace{p_{f}(\hat{s}_{i+1}|\hat{s}_{i},a_{i})}_{{\scriptsize\shortstack{simulation with\\ world model}}}}$ (5)
@@
3.4 Design of Simulative Reasoning Agent with Concept-based Latent States and Hierarchical Planning
In this subsection, we present our design of a generalizable simulative reasoning agent architecture. In particular, we provide detailed discussion on design decisions that enable robust and broad applicability across environments and tasks.